Josh is 10 years older than his brother. Next year, he will be three times as old as his brother.
a. How old was Josh when his brother was born?
b. If his brother's age is represented by a, what's Josh's age?

Answer by ptaylor(2198)

Let a=Josh's brother's age; a+1=Josh's brother's age one year from now
Then a+10=Josh's age; (a+10)+1=Josh's age one year from now
Now we are told the following:
(a+10)+1=3(a+1) dimplify
a+11=3a+3 subtract a and also 3 from both sides
a-a+11-3=3a-a+3-3 collect like terms
8=2a divide both sides by 2
a=4------------------------------Josh's brother's current age
a+10=4+10=14---------------------------Josh's current age.
Now, by inspection, we see that when Josh's brother was 4-4=0 years old, Josh was 14-4=10 years old
CK
4+10=14
14=14
14+1=3*(4+1)
15=15

Hope this helps---ptaylor